OpenStudy (anonymous):

which of the following are solutions to the equation Check all that apply. x^2-6x+9=11

OpenStudy (anonymous):

a.\[x=\sqrt{11}+3\]

OpenStudy (anonymous):

B.\[x=-\sqrt{2}+6\]

OpenStudy (anonymous):

C.\[x=2-\sqrt{6}\]

OpenStudy (anonymous):

D.\[x=-\sqrt{11}-3\]

OpenStudy (anonymous):

\[\left( x-3 \right)^{2}=11\] \[x-3=\pm \sqrt{11}\]

OpenStudy (anonymous):

F.\[\sqrt{2}+6\]

OpenStudy (anonymous):

your answer?

OpenStudy (anonymous):

E.\[x=2+\sqrt{6}\]

OpenStudy (anonymous):

dont get it

OpenStudy (anonymous):

\[x^2-6x+9=11 \]\[(x-3)^2 = 11\]Now take the square root and...

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

its A and D

OpenStudy (anonymous):

\[x-3 = \pm \sqrt{11}\]\[x = \pm \sqrt{11}+3\]What are they?

OpenStudy (anonymous):

Cannot be D... Unless D is +3 and not -3

OpenStudy (anonymous):

Did you have a typo?

OpenStudy (anonymous):

the 3 is +3

OpenStudy (anonymous):

okay! Then yes! A and D

OpenStudy (anonymous):

\[x=\sqrt{11}+3 andx=-\sqrt{11}+3\]

OpenStudy (anonymous):

thanks man hey omw to work thanks alot ima have to finish my test tomorrow

OpenStudy (anonymous):

seriously. thats exactly what i said

OpenStudy (anonymous):

lol ur going to work to

OpenStudy (anonymous):

no. "thats exactly what I said" means I told you the same thing surji said but faster/first